Circular No. 8865

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2007 Q2 (GILMORE) An apparently asteroidal object discovered by Alan C. Gilmore on CCD

images taken with the Mount John University Observatory 1-m reflector (discovery observation tabulated below), and posted on the ‘NEOCP’ webpage, has been found to show a diffuse coma and a tail 6′′–8′′ long in p.a. 320^{o} on a co-added 600-s image obtained by K. Sárneczky and L. L. Kiss on Aug. 30.6 UT with the Siding Spring 1.0-m reflector. Also, Q.-z. Ye, Guangzhou, China, writes that four stacked 120-s CCD images obtained by C.-S. Lin with the Lulin Observatory 1.02-m f/8 reflector on Aug. 27.7 UT in strong moonlight (seeing ∼ 1′′.0) show a possible weak coma of diameter ∼ 1′′.9 with a wide, short tail ∼ 3′′ long in p.a. ∼ 340^{o}–10^{o}.

2007 UT  α_{2000}  δ_{2000}  Mag.
Aug. 22.58265  22^{h}44^{m}16.^{s}31  −0^{o}46′07′′.9  19.1

The available astrometry, the following preliminary elliptical orbital elements, and an ephemeris appear on MPEC 2007-Q43.

T = 2007 Aug. 23.465 TT  ω = 162.^{o}931
e = 0.67176  Ω = 172.256 }2000.0
q = 1.83906 AU  i = 10.235
a = 5.60282 AU n^{o} = 0.074318 P = 13.26 years

COMET C/2007 Q3 (SIDING SPRING) An apparently asteroidal object discovered by D. M. Burton with the

0.5-m Uppsala Schmidt telescope in the course of the Siding Spring survey (discovery observation tabulated below), and posted on the ‘NEOCP’ webpage, has also been found to show cometary appearance by Sárneczky and Kiss, who on Aug. 30.8 UT found a 6′′ coma and 12′′-long tail in p.a. 290^{o} on a 1800-s co-added image taken with the 1.0-m reflector at Siding Spring. Also, N. Teamo and J. C. Pelle write that CCD exposures taken on Aug. 29.6 with a 0.35-m reflector at Punaauia, Tahiti, show the comet’s FWHM image to be larger than those of neighboring stars of similar brightness.

2007 UT  α_{2000}  δ_{2000}  Mag.
Aug. 25.76023  3^{h}28^{m}35.^{s}22  −46^{o}39′40′′.7  17.8

The available astrometry, preliminary parabolic orbital elements (T = 2009 Oct. 19.198 TT, q = 2.31933 AU, ω = 0^{o}.675, Ω = 151^{o}.078, i = 66^{o}.275, equinox 2000.0), and an ephemeris appear on MPEC 2007-Q44.

2007 August 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green